### v1

# Theory uncertainties of the irreducible background to VBF Higgs production

Abstract
Higgs boson production through gluon fusion in association with two jets is an irreducible background to Higgs boson production through vector boson fusion, one of the most important channels for analyzing and understanding the Higgs boson properties at the Large Hadron Collider. Despite a range of available simulation tools, precise predictions for the corresponding final states are notoriously hard to achieve. Using state-of-the-art fixed-order calculations as the baseline for a comparison, we perform a detailed study of similarities and differences in existing event generators. We provide consistent setups for the simulations that can be used to obtain identical parametric precision in various programs used by experiments. We find that NLO calculations for the two-jet final state are essential to achieve reliable predictions.
I Introduction
The discovery of the Higgs boson Aad et al. (2012); Chatrchyan et al. (2012) by the ATLAS and CMS experiments at the Large Hadron Collider (LHC) has ushered in a new era in elementary particle physics. The so far excellent agreement of the boson’s properties with Standard Model expectations leaves as many questions open as it answers, and calls for a more detailed investigation of this unique particle. One of the most promising Higgs production modes for in-depth studies is weak vector boson fusion, conventionally referred to as VBF. This process has been analyzed by both ATLAS and CMS for a number of Higgs boson decay channels at a number of center-of-mass energies Aad et al. (2022); Tumasyan et al. (2022). High-statistics measurements of VBF allow for a precise determination of the couplings of the Higgs boson to and bosons, as well as improved limits in searches for possible new physics. The presence of two (or more) jets in the final state, characterized by a large rapidity separation and/or a large dijet invariant mass, can be used to enhance the VBF sub-process over the dominant production mechanism, which is gluon-gluon fusion (ggF). However, the final-state configurations in gluon fusion with two (or more) jets can still be identical to the one in VBF, turning ggF from a signal into an irreducible background. The full potential of the LHC for VBF Higgs boson measurements can therefore only be realized, once gluon fusion Higgs production with two additional jets is theoretically understood to the highest achievable precision.
Matrix element plus parton shower matching Frixione and Webber (2002); Nason (2004); Frixione et al. (2007); Alioli et al. (2010); Hoche et al. (2011); Höche et al. (2012); Alwall et al. (2014); Jadach et al. (2015); Nason and Salam (2022) and merging André and Sjöstrand (1998); Catani et al. (2001); Lönnblad (2002); Mangano et al. (2002); Krauss (2002); Lavesson and Lönnblad (2008a); Höche et al. (2009); Hamilton et al. (2009); Lönnblad and Prestel (2012, 2013a); Plätzer (2013); Höche et al. (2019); Lavesson and Lönnblad (2008b); Hoeche et al. (2013); Frederix and Frixione (2012); Lönnblad and Prestel (2013b); Hoeche et al. (2014) techniques are an essential theory tool to achieve this goal, and are used in the extraction of the Higgs boson cross sections from experimental measurements. They are embedded in complete frameworks for the production of the particle-level final state provided by the general-purpose LHC event generators, Herwig, Pythia, and Sherpa Buckley et al. (2011); Campbell et al. (2024). Many other theory tools make use of the event generators to provide the implementation of resummation and/or hadronization in terms of parton showers and fragmentation models. The interplay between perturbative and non-perturbative QCD is a key aspect of Monte Carlo event generation at the LHC. Simulations typically assume a scale hierarchy, where the hardest interactions of the event are described by perturbatively computed matrix elements, while the parton shower adds softer radiation using resummed perturbation theory. At some point during the QCD evolution, a scale is reached where the transition to the non-perturbative regime must be implemented. The simulation of this hadronization process relies on physically motivated models that are tuned to experimental data from current and past experiments, notably data from LEP, HERA and the LHC. As these data sets are the same for all event generators, the different hadronization models usually predict fairly similar results, and their uncertainties tend to be of a similar size Höche et al. (2012); Alcaraz Maestre et al. (2012), although in cases where particle spectra are resolved in detail, significant differences can still occur Carli et al. (2010); Divisova et al. (2025). On the other hand, one often observes discrepancies in the simulation of perturbative QCD effects. A classical example would be if the parton shower settings involved a different form of the strong running coupling than the hard matrix elements, or if the value of the running coupling differed at the reference scale. Such mismatches are due to the attempt to “tune away” effects that can arise from missing higher-order corrections. Inconsistent tunes, however, are not predictions based on solid QCD perturbation theory, and the corresponding simulation results can no longer be relied upon.
In a previous publication Buckley et al. (2021), some of us focused on a consistent way to simulate VBF Higgs boson production at both fixed-order (LO, NLO, NNLO) and fixed-order plus parton shower, finding agreement in both the central predictions and their uncertainties at the level expected of the perturbative QCD result. These calculations are in line with the observations of a range of other studies Jäger et al. (2014, 2020); Ballestrero et al. (2018); Höche et al. (2022); Bittrich et al. (2022); Chen et al. (2022a); van Beekveld and Ferrario Ravasio (2024). In the present paper, we will explore the gluon-gluon fusion background, using a variety of tools in the same common setup as required to obtain theoretically consistent results.
The analyses performed here match those used by ATLAS and CMS as much as possible. We will show that the nominal setups for the different programs lead to similar predictions for the kinematic distributions crucial to the determination of the VBF Higgs signal. As a result, the uncertainties quoted for the ggF backgrounds by ATLAS and CMS may be overstated. We will compare predictions from approaches that are inclusive in the number of jets (; such as from Sherpa MEPS@NLO) to those from more exclusive approaches, where only the final 2 jet state is produced (such as Sherpa S-MC@NLO and Powheg Box + Pythia/ Powheg Box + Herwig ). We also calculate the MiNNLO prediction for gluon-fusion (ggF) Higgs boson production, in which the leading jet spectrum is computed at NLO precision, while the second jet is predicted at LO. This is the framework which is currently used by the ATLAS and CMS experiments. We compare results at the parton shower level and the hadron level and discuss similarities and differences. We disentangle uncertainty estimates as much as possible, in order to provide a better understanding of their sources. We also test the impact of the merging scheme and scale. We provide a Rivet routine that can be directly used by the experiments to confirm the correct use of each generator. Although the ATLAS and CMS determinations of the VBF Higgs cross section are not strictly cut-based, we will also discuss a cutflow analysis, indicating the results of successive phase-space restrictions that are used in VBF Higgs analyses, as well as providing comparisons of what has been termed loose and tight VBF selection cuts.
This manuscript is structured as follows: Section II introduces the event generators we use for the study, and explains the setup of the perturbative calculation and the non-perturbative tunes. It also lists the setups in which the event generators are typically employed by the experiments. In Sec. III we compare the results from those generators to the predictions typically obtained from the current LHC simulations. Section IV summarizes our findings on uncertainties, while Secs. V and VI present some additional phenomenological results. Section VII contains our conclusions and an outlook.
II Input parameters and Computational Tools
To provide meaningful theoretical guidelines for experimental analyses and simplified template cross-sections (STXS), we adopt the same jet definition as in ATLAS measurements Gamel et al. (2019); ATL (2019) and study differential observables using fiducial bins suggested by STXS stage 1.1 Berger et al. (2019). We focus on on-shell Higgs bosons of GeV, produced in the ggF channel with a center-of-mass energy of 13.6 TeV. The Higgs vacuum expectation value is . Theoretical uncertainties are estimated by varying QCD renormalization () and factorization () scales independently by a factor of two around the central scale of , where
We eliminate the two antipodal combinations and . The parton distribution functions (PDFs) are defined by the central group of PDF4LHC21_40 NNLO PDFs Ball et al. (2022). The computational tools used in this study are introduced in the following.
II.1 NNLOJet
NNLOJet is a parton level event generator specializing in the computation of NNLO QCD corrections Huss et al. (2025). The antenna subtraction formalism Gehrmann-De Ridder et al. (2005a, b); Daleo et al. (2007, 2010); Gehrmann-De Ridder et al. (2012); Currie et al. (2013) is used to regulate IR divergences at each stage of the fixed-order calculations and to provide fully-differential results. For Higgs production in the gluon-fusion channel, NNLOJet provides predictions at NNLO Chen et al. (2015, 2016). This implies that for observables with at least one (), two () and three jets (), the corresponding fixed order QCD accuracy is NNLO, NLO and LO, respectively. In this study, the calculation is performed in the Higgs effective field theory (HEFT) with infinitely large top-quark mass. The relevant Standard Model (SM) top-quark mass effects were studied for at NLO with full theory approximation on the two-loop matrix elements finding comparable NLO/LO ratios between HEFT and SM for differential observables Chen et al. (2022b), justifying the treatment of top-quark mass in the rest of this study. More details on this can be found in App. A.
II.2 POWHEG Box
The Powheg Box Frixione et al. (2007); Alioli et al. (2010); Ježo and Nason (2015) is an event generator which produces Les Houches events Boos et al. (2001); Alioli et al. (2014) with NLO accuracy that can be matched to parton showers according to the POWHEG method Nason (2004). In Powheg Box, the following main simulation modes for ggF are available and used in this study:
boson production at NNLO QCD Monni et al. (2020) implemented in the MiNNLO framework (first achieved by reweighting in association with one extra jet in the MiNLO framework in Hamilton et al. (2013a)). In this setup, the Higgs boson production without additional jets () is NNLO accurate, is NLO accurate and is LO accurate.
at NLO QCD Campbell et al. (2012) implemented in the plain Powheg Box, with and not provided and with at NLO precision.
We note that the MiNNLO-based prediction for the 2-jet final state is leading-order accurate. Next-to-leading order precision is achieved by using the simulation, which requires generation cuts or Born suppression (see Sec. IV.2 for details). While Born suppression works well at the parton level, full event generation can fail if the simulation of multiple parton interactions is allowed to generate jets at transverse momenta larger than that of the hard process. Events with low transverse momentum jets in the hard process, which have large (and in principle unbounded) weights, can then be elevated to the 2-jet final state, which will result in distributions with seemingly random statistical outlier events. We therefore decide to define the upper scale for both parton-shower and MPI evolution as the value given by the SCALUP parameter. This technique is not available in Herwig, therefore comparisons including Powheg Box + Herwig will be limited to the parton level.
We have cross-checked the Powheg Box NLO predictions at fixed order against NNLOJet, Herwig and Sherpa. Note that a sample with all three multiplicities (, and ) provided at NLO accuracy can be obtained with an extension of the MiNLO method described in Ref. Frederix and Hamilton (2016), but the corresponding code is currently not publicly available.
II.3 Pythia
Pythia Sjöstrand et al. (2006); Bierlich et al. (2022) is a multi-purpose particle-level event generator with a historically strong focus on soft physics. The non-perturbative evolution from colored partons to color-singlet hadrons is modeled using its distinctive feature, the string-fragmentation model Andersson et al. (1983); Sjöstrand (1984). In the present version 8.3, Pythia implements the Vincia antenna shower Brooks et al. (2020) in addition to the default -ordered “simple shower” Sjöstrand and Skands (2005)111In versions prior to 8.316, also the Dire dipole-like parton shower Höche and Prestel (2015) was available.. For the simple shower, a dipole-recoil option is available, which replaces the separate incoherent initial-final and final-initial parton evolution by a common, coherent initial-final evolution Cabouat and Sjöstrand (2018). For each of the parton-shower algorithms, multi-parton interactions are fully interleaved with the shower evolution Sjöstrand and Skands (2005). In this study, we use Pythia 8.315.
Pythia 8.3 implements a series of processes at LO. Both the VBF Higgs-production process and the gluon-fusion Higgs production process in the heavy-top limit are available at LO accuracy. In case of the latter, the simple shower algorithm also incorporates LO matrix-element corrections for the first emission, corresponding to the configuration. For matching and merging, Pythia relies on Les Houches events Boos et al. (2001); Alioli et al. (2014) from external programs such as MadGraph5_aMC@NLO Alwall et al. (2014), Powheg Box, or Sherpa. Multi-jet merging is implemented in the CKKW-L approach Lönnblad (2002); Lönnblad and Prestel (2012). For NLO matching with the Powheg Box, dedicated PowhegHooks are available for all internal parton-shower algorithms to account for the mismatch between the evolution variable in Powheg Box and the shower evolution variable in Pythia Corke and Sjöstrand (2010). NLO matching with MadGraph5_aMC@NLO is only available with the default simple shower with (non-default) global longitudinal recoil in final-state branchings, which was found to be problematic for certain simulations and will therefore not be considered in this study.
II.4 Herwig
Herwig Marchesini and Webber (1984, 1988); Bähr et al. (2008) is a general-purpose event generator with a historically strong focus on perturbative QCD and its connection to analytic resummation. Its distinctive feature is an angular-ordered parton shower (see Refs. Gieseke et al. (2003); Bewick et al. (2020) for a modern reformulation of such a shower) and a cluster hadronization model. Its latest version, Herwig 7.3 Bellm et al. (2020); Bewick et al. (2024), also implements a dipole shower Platzer and Gieseke (2011), as well as an electroweak parton shower Masouminia and Richardson (2022). Herwig 7.3 allows for flexible NLO matching and merging capabilities. Thanks to its Matchbox module Plätzer and Gieseke (2012), both Powheg Nason (2004) and MC@NLO Frixione and Webber (2002) matching are offered, as well as unitarized merging Bellm et al. (2018) (just for the dipole shower). It also supports showering of Les Houches events generated either by the Powheg Box or by MadGraph5_aMC@NLO (matching to MadGraph5_aMC@NLO is supported only for the angular-ordered shower).
In this article, we use the angular-ordered shower to process Les Houches events generated by the Powheg Box. In particular, we use the powhegHerwig plugin developed in Ref. Ferrario Ravasio and Oleari (2023), which is largely based on the Les Houches interface provided by the ThePEG Lonnblad (2009). Independently, Herwig’s fixed-order calculation of the cross section and distributions has been checked against NNLOJet, Powheg Box and Sherpa.
II.5 Sherpa
Sherpa Gleisberg et al. (2004, 2009); Bothmann et al. (2019, 2024) is a general-purpose event generator with a strong focus on perturbative physics, in particular multi-jet merging methods. We use version 3.0 of the program for this study. The matching of fixed-order NLO corrections to the parton shower is performed in the S-MC@NLO approach Höche et al. (2012, 2013), while NLO computations are performed using Amegic Krauss et al. (2002); Gleisberg and Krauss (2008), Comix Gleisberg and Höche (2008), and MCFM Campbell and Ellis (1999); Campbell et al. (2011, 2015); Campbell and Neumann (2019); Campbell et al. (2021). We have cross-checked the virtual corrections against OpenLoops Cascioli et al. (2012); Buccioni et al. (2018, 2019) and compared the complete NLO predictions against NNLOJet, Herwig and Powheg Box. We use a modified version of the Catani-Seymour dipole shower algorithm (CSS) Schumann and Krauss (2008), which was cross-checked against the Dire parton shower Höche and Prestel (2015) (for S-MC@NLO only). The latter has been deprecated in Sherpa, and therefore we only present predictions using the default CSS algorithm. The NLO multi-jet merging is based on Gehrmann et al. (2013); Hoeche et al. (2013). Hadronization in Sherpa is provided by a cluster fragmentation model Winter et al. (2004); Chahal and Krauss (2022), and multiple parton scattering as a model of the underlying event is simulated according to Alekhin et al. (2006a, b).
III Comparison to typical LHC simulations
Before performing a detailed analysis of theoretical uncertainties, we compare our new predictions to those obtained from a setup routinely employed by the experimental collaborations. This provides the context for the discussion of the true uncertainties in Sec. IV. The event generator conventionally used by ATLAS and CMS for both VBF Higgs boson production and the ggF Higgs production background was NNLOPS Hamilton et al. (2013a, 2015). It has now largely been replaced by the more modern MiNNLO generator Monni et al. (2020); Niggetiedt and Wiesemann (2024), which gives formally equivalent results while being substantially more efficient, due to the elimination of the reweighting step. Both NNLOPS and MiNNLO are usually interfaced with the Herwig and Pythia Monte-Carlo programs. The estimated uncertainty in the ggF Higgs boson background is dominated by the observed differences between the predictions of NNLOPS+Pythia and NNLOPS+Herwig, which arise from both perturbative and non-perturbative QCD effects. A key point is that these are the differences observed by using the simulations “out of the box” (albeit with tuned non-perturbative parameters from each experiment), and not differences resulting from true discrepancies in the physics models of the event generators. In this section, we will use the ATLAS and CMS setups of MiNNLO for comparison to the ones we recommend as a new baseline to reach the highest theoretical precision in jets from ggF. Note in particular that the use of NNLOPS to simulate the ggF background to VBF Higgs production implies that the final state is predicted at leading order accuracy only, which is below the current state of the art. We also stress that in order to combine predictions from a Powheg Box generator with an angular-ordered shower, one would need to include truncated showers to fill the large-angle region Nason (2004). This also applies even when Powheg Box events are generated in accordance with the NNLOPS or MiNNLO methods, where now one has to imagine the two hardest emissions having been pre-determined. Truncated showers are implemented in the internal Herwig implementation of POWHEG, but not if one showers (NLO or NNLO) events generated with external tools. One can argue that at leading color, simply vetoing radiation with a transverse momentum larger then the (softest) Powheg Box emission is sufficient to fill the large-angle region while retaining leading logarithmic accuracy (see e.g. Appendix D of Ref. Ferrario Ravasio et al. (2018)). The same reasoning applies if one considers two emissions as in the case of NNLOPS/MiNNLO. In the case of POWHEG matching, the effect of truncated showers has been studied e.g. in Ref. Hamilton et al. (2008), and found to be small, but such a study is missing in the case of two hard emissions. We give more details on the precise Monte-Carlo simulations used in the experiments in App. C.
III.1 Observable definitions
We will discuss the following four observables, which are sensitive to the characteristics of the Higgs-plus-two-jet final state:
the transverse momentum of the Higgs boson, ;
the invariant mass of the two leading jets, ;
the rapidity separation of the two leading jets, ;
the azimuthal separation of the two leading jets, .
Additional data are collected on https://doi.org/10.5281/zenodo.17015614, together with the Rivet Buckley et al. (2013); Bierlich et al. (2020, 2024) analysis routine. The Higgs boson is required to be within the central region of the detectors () in association with at least two accompanying jets. The jets are clustered using the FastJet implementation of the anti- algorithm Cacciari et al. (2012), and must satisfy the following cuts:
We primarily consider jets with a radius of 0.4, but also provide results for radii of 0.3, 0.6 and 0.8, to quantify the impact on selected differential observables.222Our study of the VBF Higgs boson production process in Ref. Buckley et al. (2021) examined jet -values from 0.3 to 1.0, and indicated a strong -dependence due to the kinematic requirements on the two jets. The R-dependence increased upon the imposition of the VBF phase space cuts, so it is important to check if the same effects occur for the background. Note that the two leading jets in VBF Higgs boson production are primarily quark jets, while in ggF they are primarily gluon jets. This may led to some differences in the -dependence. Jets are ranked by the magnitude of their transverse momentum.
To distinguish between normalization and shape differences of the predictions, all results will be normalized to the inclusive cross section for the final state, defined at the particle level. This means in particular that a normalization difference can occur through the different behavior of the parton shower simulation, for example if the parton shower shifts a jet above/below the cut by means of additional radiation and/or recoil effects. The cross section is inferred from a relatively inclusive distribution, which we choose to be the Higgs boson rapidity distribution in the two jet final state.
III.2 Results of the comparison
We are now ready to compare results from the MiNNLO simulation as used by ATLAS and CMS to those from the tools introduced in Sec. II. The MiNNLO calculation is matched to the default Pythia parton shower (“simple shower”), both with and without the “dipole-recoil” option, and to the Herwig angular-ordered parton shower. The results are compared to a MEPS@NLO simulation from Sherpa, including the , , and final states at NLO precision, and up to at LO accuracy. This sample will be dubbed MEPS@NLO in the following. All predictions are at parton level, without multi-parton interactions and non-perturbative corrections. We also use a fixed-order NNLO calculation from NNLOJet as a reference. As a reminder, MiNNLO is NNLO accurate for the inclusive Higgs boson cross section, NLO accurate for the 1-jet final state, and LO accurate for the 2-jet final state.
Figure 1 shows these comparisons. The middle panels of each sub-figure display the ratio to the inclusive fixed-order calculation from NNLOJet, while the bottom panels compare the event-generator predictions, normalized to the MiNNLO + Pythia result. We find approximate agreement of the rate in the Higgs rapidity distribution of the selection (not shown here) between NNLOJet and MiNNLO + Pythia, whereas the Sherpa MEPS@NLO sample requires a scale factor of , and the MiNNLO + Herwig sample a scale factor of . In the case of the Sherpa prediction, the scale factor is consistent with the NNLO to NLO ratio of the inclusive Higgs boson production cross section Anastasiou et al. (2016), which is needed to bring the inclusive jet rates in the MEPS@NLO result into approximate agreement with the prediction from NNLOJet.
This can be understood better with the help of Fig. 2, which shows the leading jet transverse momentum spectra in the inclusive and selection. Formally, the MiNNLO simulation models the first of these spectra at NLO precision, and the second at LO precision. The MEPS@NLO simulation should predict both spectra at NLO precision, but with added Sudakov suppression effects. The NNLOJet reference provides NNLO precision for the first spectrum, and NLO precision for the second. Figure 2 shows that the scaled MEPS@NLO result has the correct shape over most of the fiducial phase space, and in the medium hard region in particular. The MiNNLO predictions on the other hand agree neither with the leading jet spectrum in the selection, nor with the leading jet spectrum in the selection. In particular, we observe a significant slope difference in the medium hard region of the leading jet transverse momentum in the selection, where the MiNNLO result is formally NLO accurate. We therefore conclude that the agreement in overall rate between the MiNNLO + Pythia result and the NNLOJet prediction is rather accidental. We will discuss this point in more detail in Sec. IV.4, and in particular when comparing the distributions in Fig. 9.
The rapidity separation and the invariant mass distribution of the two leading jets, shown in Figure 1, agree well between the Sherpa, Pythia, and Herwig setups. At low invariant masses, , and small rapidity separations, , all results also agree well with NNLOJet, while they predict a 20-30% increased rate in the complementary region. The distributions of the azimuthal separation generally shows good agreement between the different predictions in the region of small and intermediate separations, where they lie within a 10% envelope of each other. At large jet separations, , the MiNNLO + Pythia prediction is larger than all other predictions by about 30%, including the MiNNLO + Pythia results using the dipole-recoil option. We will discuss this in more detail in Fig. 9. The Higgs transverse-momentum distribution shows the largest shape differences between the different simulations. At low transverse momentum, , both MiNNLO + Pythia results and MiNNLO + Herwig exceed the NNLOJet result by about 20-30%, with close agreement between the MiNNLO predictions obtained with Herwig and with the Pythia dipole-recoil shower. The Sherpa MEPS@NLO prediction is about 20% lower than the NNLOJet result. Starting from , both MiNNLO + Pythia predictions agree very well with NNLOJet, whereas the Sherpa MEPS@NLO and MiNNLO + Herwig predictions create an envelope around the fixed-order result, with Sherpa MEPS@NLO being larger by about 10% and MiNNLO + Herwig being smaller by about the same amount. As can be inferred from the middle panes in all four figures, the event-generator predictions broadly lie within the seven-point scale variation uncertainty band of the fixed-order calculation, with some exceptions towards larger values of the observables.
Based on the observations in Fig. 1, one may be tempted to conclude that the modeling of the irreducible background to VBF Higgs production suffers from large theory uncertainties. This is in fact not the case. The event-generator predictions compared so far have rather different formal accuracy in the fiducial phase space, and therefore provide a different level of theoretical control over the observables of interest. Furthermore, as already discussed earlier, while in the case of NLO matching with an angular-ordered shower, truncated showers are known to yield a very small effect, an in-depth study is missing for the case in which the two hardest emissions have already been generated. Most of these effects can be mitigated once the final state is modeled consistently at NLO accuracy. This will be discussed in the next section.
IV Assessment of theory uncertainties
In the following, we will use a consistent baseline for the hard, perturbative calculation – at NLO precision – and assess the influence of different theory variations on the size of the theoretical uncertainties in the fiducial phase space. Throughout the discussion, we will focus on the same four observables as in Fig. 1, with additional results available at https://doi.org/10.5281/zenodo.17015614. In Sec. IV.1, we compare parton-shower variations in NLO-matched distributions at the parton level. Sec. IV.2 focuses on matching variations in Powheg Box, while the influence of dedicated Pythia tunes and different hadronization models in Sherpa is discussed in Sec. IV.3. A study of the impact of the Higgs effective theory approximation can be found in App. A and a study of the jet-radius dependence can be found in App. B.
IV.1 Parton-shower variations
In this section, we compare the influence of different parton-shower algorithms, matched to a NLO calculation. We consider a total of four different showers in Sherpa, Pythia, and Herwig. In Sherpa, we use the Catani-Seymour shower matched to the NLO calculation using the S-MC@NLO scheme; in Pythia, we employ the default -ordered “simple” shower as well as the simple shower with dipole recoil, both matched to the NLO calculation using Powheg Box; in Herwig, we use the angular-ordered shower matched to the NLO calculation in the POWHEG scheme using Powheg Box. A discussion of uncertainties pertaining to the matching scheme is given in Sec. IV.2. In all cases, we consider parton-level predictions without modeling the underlying event and we do not include hadronization effects. The impact of these corrections is addressed in Sec. IV.3.
As before, we scale all distributions to the NNLOJet rate based on the Higgs rapidity spectrum, yielding scale factors of in the case of Sherpa, for both Pythia showers, and for Herwig. The resulting parton-shower variations are shown in Fig. 3, including the NNLOJet NNLO prediction as a baseline. Besides the main distributions, we also show two ratio panels, with the first (middle panel) comparing the matched predictions to the NNLOJet fixed-order reference, and the second (bottom panel) showing the difference between the matched parton-shower predictions, normalized to the Powheg Box + Pythia result. All predictions in Fig. 3 acquire formal NLO accuracy in configurations.
As can be seen from Fig. 3, we find excellent, percent-level, agreement between the four matched parton-shower algorithms for the , , and distributions. For the distribution, both Pythia showers agree very well with Sherpa, while the Herwig angular-ordered shower differs from these at large transverse momentum, . In all cases, the matched parton-shower results are within the scale uncertainties of the fixed-order calculation, although the agreement with NNLOJet is considerably better for and than for and , with the latter showing the largest differences of up to 20%.
IV.2 Matching-scheme variations
Having observed very good agreement between different parton-shower prediction and the fixed-order calculation, we now investigate ambiguities in the matching scheme. Since different matched parton-shower predictions generally agreed very well, we focus on Powheg Box + Pythia, where we expect the largest differences arising from ambiguities in the multiplicative POWHEG approach and the interfacing to an external shower generator. Specifically, we study the influence of the following Powheg Box event-generation parameters:
the cut-off scale of the first emission (handled by Powheg Box), ptsqmin;
the scale choice in real contributions, btlscale, namely whether to use the real kinematics of the corresponding Born projection for the computation of the renormalization and factorization scale (see Sec. 3 of Ref. Ferrario Ravasio and Oleari (2023));
hdamp, i.e. the transverse-momentum scale used to separate regular real events from soft ones (see Sec. 5 of Ref. Alioli et al. (2010) for further details).
In addition, there is some ambiguity in how the scale mismatch between the first emission in Powheg Box and Pythia is treated Corke and Sjöstrand (2009). Natively, Pythia uses PowhegHooks to veto emissions harder than the first emission of Powheg Box. By default, however, the process in Powheg Box regulates the process by a procedure known as Born suppression, in which infrared divergent configurations are suppressed. This approach differs from the usual generation cuts in that events with singular configurations are not entirely excluded from the event sample but enter the sample with a reduced (or vanishing) weight. Typically, these singular events are excluded from the analysis by the fiducial cuts, which separate out the topology. However, we found that in both Pythia and Herwig this approach suffers from deteriorating statistical precision, as jets created by multi-parton interactions can migrate (previously suppressed) infrared-divergent events into the fiducial phase space (see also Sec. II.2). By default, we have therefore decided against using PowhegHooks in Pythia for our main study, meaning that both the parton-shower and MPI evolution is started at the Powheg Box SCALUP scale without further vetoes. A similar approach does not seem viable in the underlying-event simulation in Herwig, where the MPI starting scale cannot be altered. This excludes full event simulations, including hadronization and MPI, with Powheg Box + Herwig from our study. We hope that it will be possible to implement this feature in Herwig in the future.
In order to gauge the importance of vetoed showers in our predictions, we have implemented generation cuts in the Powheg Box process and compare the following event-generation setups:
default Born suppression without PowhegHooks: the importance sampling is altered so to have many events at large , and few events with large weight at small ;
tighter generation cuts without PowhegHooks;
tighter generation cuts with PowhegHooks and POWHEG:pThard = 0;
tighter generation cuts with PowhegHooks and POWHEG:pThard = 2.
In this context, the POWHEG:pThard setting defines how the veto scale is calculated in the PowhegHooks, with POWHEG:pThard = 0 corresponding to the scale of the Powheg Box emission, SCALUP, and POWHEG:pThard = 2 amounts to recalculating the veto scale as the minimal across the entire event. With properly adjusted parton-shower and MPI starting scales, the first two options should fully agree. The third option provides a proxy for the importance of vetoed showers in our setup, while the last option corresponds to an alternative veto-scale choice in PowhegHooks, which was recently adopted in the ATLAS experiment to quantify the evolution-scale mismatch between Powheg Box and Pythia The ATLAS Collaboration (2023). It should be noted that sizable differences between the latter two options have been observed for the VBF process in Höche et al. (2022).
A comparison of the different matching options is shown in Fig. 4, where the lower panel compares the various settings to the default approach using Born suppression without PowhegHooks. In case of the btlscale = 0 (i.e. using the real kinematics to compute the scales entering the real cross section) and hdamp = 125 predictions, we again apply scaling factors of and , respectively, based on the Higgs rapidity distribution. We observe excellent agreement in the shape of all distributions, with differences well below the percent-level. The largest differences are found with the hdamp = 62.5 predictions, which differs from the other predictions by up to 10% in the tails of the , , and distributions, and by about 5% in the region of small Higgs transverse momentum, . This emphasizes that within the freedom of the Powheg Box matching implementation, consistent results are obtained as long as the parameters do not take on extreme values. We find that the btlscale setting has a small impact on the shape of distributions, while a small value of hdamp can cause an important shape difference at large /. This is due to the fact that, in the POWHEG method, for relatively large transverse-momentum radiation, the predicted rate is effectively equivalent to the real cross section multiplied by the ratio of the inclusive NLO/LO cross section, which is positive for the process at hand. When hdamp is used, radiation harder than this parameter is treated separately and effectively does not include this factor.
IV.3 Tune variations
We next turn to study uncertainties arising from differences in the modeling of non-perturbative and underlying-event effects. To this end, we consider the MC@NLO prediction obtained with Sherpa in conjunction with either Sherpa’s cluster-fragmentation Chahal and Krauss (2022) or the string-fragmentation model in Pythia, interfaced at runtime. In addition, we consider three different tunes in the Powheg Box + Pythia setup with the default shower (without dipole recoil): the default Monash2013 tune in Pythia Skands et al. (2014), the ATLAS AZNLO tune Aad et al. (2014), and the CMS CP5 tune Sirunyan et al. (2020).
Fig. 5 shows the comparison of predictions for the same observables investigates earlier. The differences originating in different hadronization and underlying event models, as well as the different tunes, are within 5%, with the largest discrepancy observed for the CMS C5 tune compared to the ATLAS AZNLO tune. We note that Herwig predictions cannot be compared for this setup, because the simulation of the underlying event does not allow a restriction of the hardness of the secondary scatterings.
IV.4 Comparison of inclusive and exclusive setups
Finally, we contrast inclusive and exclusive predictions in the fiducial phase space. Specifically, we consider the Sherpa MEPS@NLO results with the MiNNLO + Pythia and MiNNLO + Herwig predictions as prototypes of inclusive results and compare them to the NLO predictions obtained with Sherpa MC@NLO, Powheg Box + Pythia, and Powheg Box + Herwig. Caution is advised in interpreting this comparison, as the MiNNLO predictions only achieve formal LO accuracy for configurations (and the logarithmic accuracy in case of MiNNLO + Herwig might not even be leading logarithmic), while all other setups, including the Sherpa MEPS@NLO prediction, are NLO precise with leading logarithmic accuracy on the parton shower side.
Predictions for the , , and are shown in Figs. 6-9. There is some difference between modeling higher jet multiplicities at leading order (Sherpa MEPS@NLO) and treating them as part of the inclusive calculation (Sherpa MC@NLO, Powheg Box + Pythia, Powheg Box + Herwig), in particular for observables that are sensitive to additional hard jets, such as . For variables such as , and , differences are smaller.
Perhaps the most striking feature is observed in the comparison of , shown in Fig. 9. There is a marked shape difference between MiNNLO and MEPS@NLO in the region around , a region important for studies of CP-violation, for both VBF Higgs and ggF production. This difference is amplified at the parton level. Moreover, there is a dramatic shape difference between MiNNLO + Herwig and MiNNLO + Pythia, with MiNNLO + Herwig indicating a similar behavior around as the NLO matched predictions. As a consequence of these observations, we conclude that the MiNNLO simulation has significant deficiencies when it comes to predicting the azimuthal correlations between the leading jets. This adds to the problems of describing the leading jet transverse momentum spectra observed in Fig. 2. On the other hand, we observe that all calculations that are NLO precise for the final state are consistent in their predictions of the spectrum. This includes in particular the MC@NLO predictions from Sherpa, the Powheg Box + Pythia and Powheg Box + Herwig predictions, and the inclusive MEPS@NLO predictions from Sherpa. We have verified that the MEPS@NLO results also agree with LO merged results from Sherpa, and that the latter are stable with respect to the number of jets included in the merging, as long as a third jet is included. Our recommendation is therefore to rely on any one of those predictions by default, and to gauge the uncertainty of the result by varying among the NLO accurate setups presented in this study.
V The transverse momentum of the Higgs-plus-dijet system
In this section we discuss the transverse momentum spectrum of the Higgs boson and the two leading jets. This observable is predicted at leading order precision in the NLO accurate simulations of the final state employed in this paper. It probes both the real radiation pattern in the parton-shower approximation, and the effect of the choice of renormalization and factorization scale. It is therefore an excellent gauge of the residual differences between the various event generator predictions at the parton level.
Figure 10 (left) shows the results for NLO matched calculations. The agreement between most of the predictions is at the level of 10%, which is well within the range expected from renormalization scale variations for this complex final state. The largest difference is observed between the “simple shower” result from Powheg Box + Pythia and the prediction from Powheg Box + Herwig. In this case, the deviation is about 20%, which can still be considered a good agreement, given the complexity of the Born process.
If an experimental analysis depends significantly on the precise prediction given by an event generator for the transverse momentum spectrum of the system, an improved simulation would be required for this observable. In particular, the result should then be stabilized by including the final state at NLO precision through multi-jet merging. While this would in principle be possible with the tools employed in this publication, it would require an outsized investment into computing time, and we therefore postpone the corresponding study to a future publication.
VI Cross section in cut based analyses
In previous sections, we have been comparing relatively inclusive distributions generated by the different Monte Carlos/fixed order programs. The determination of a VBF signal/background region relies on cuts on multiple variables, so in this section we describe a cutflow that examines different kinematic regions, some of which are related to the STXS analyses.
In Fig. 11, we show the cross sections for a number of predictions which are NLO accurate in the two jet phase space, compared to the prediction from MiNNLO. Bin 0 shows the inclusive two-jet prediction, satisfying the rapidity and transverse-momentum cuts.333The Sherpa MC@NLO prediction in Fig. 11 left does not exactly reproduce the inclusive rate of the parton-level result in Fig. 11 right (Bin 0 of the histogram). This is due to the generation cuts and the very loose jet requirements, which result in a larger effect from jet production through multiple parton scatterings than in Sherpa MEPS@NLO and Powheg Box+Pythia. A more inclusive simulation (such as Sherpa MEPS@NLO) is needed if one wishes to resolve the final state in this region of the phase space. The left panel shows the results at the hadron level with MPI, and the right panel for different NLO+PS setups at the parton level. All predictions agree well, as expected given the normalization conditions described earlier. In bins 1-6, minimum cuts are placed on the dijet mass of 60, 120, 350, 700, 1000 and 1500 GeV. There is reasonable agreement with MiNNLO (on the order of 10%), but MiNNLO predicts higher invariant mass final states than the other simulations. In bins 7-11, dijet rapidity separation cuts of 1, 2.0, 2.5, 3.0, 3.5 are applied. There is good agreement among the predictions, although Sherpa MC@NLO predicts smaller rapidity separations than the other simulations.
Bins 12-17 examine the impact of cuts on the combined Higgs plus dijet transverse momentum, of 5, 10, 15, 20, 25, 30 GeV. The 25 GeV cuts is the default used in the simplified template cross section analysis. There is good agreement for the 5 GeV cut, as expected since this retains the bulk of the cross section. The MiNNLO prediction shows a different slope than the other predictions, leading to a more than 10% lower cross section for MiNNLO for greater than 25 GeV. Bins 18-21 examine the impact of cuts on the Higgs boson transverse momentum of 60, 120, 200, and 400 GeV. The differences can reach up to 15%.
Finally, in bin 22 we impose a loose VBF cut ( GeV and ), and in bin 23, we impose a tight VBF cut ( GeV and ). The Monte-Carlo predictions agree to within 10% on the rate for these two bins.
VII Conclusions
In this study, we have compared theory predictions for the irreducible background to Higgs production in vector boson fusion from gluon-gluon fusion. We began with the prevailing simulation tools used by in ATLAS and CMS, which are based on NNLO precise calculations for the inclusive Higgs boson production process, matched to a parton shower. These calculations achieve LO precision for the final states of interest in the VBF measurements and are therefore below the state of the art. In addition, differences in excess of 20% have previously been observed between different modes of running these simulations. We have pinpointed some of these differences, which have the potential to affect the determination of the CP properties of the Higgs boson.
We have also constructed a theoretically consistent set of simulation setups for the general-purpose event generators Herwig, Pythia, and Sherpa, matched to either the Powheg Box or the Sherpa internal S-MC@NLO simulation of production at NLO precision. These tools achieve formal NLO precision for the ggF induced background to VBF measurements and have been validated against NNLOJet. Comparisons among these tools show consistent predictions, at the level of 10% or less, which have been tested against variations of the matching scheme, the parton shower implementation, the hadronization model, the simulation of the underlying event, and the non-perturbative tuning parameters. In addition, we have compared the matched predictions to leading- and next-to-leading order multi-jet merged results, which provide an inclusive prediction that (in the case of NLO merging) is able to achieve NLO precision for the final state as well.
Overall, we find consistent behavior and agreement of all NLO matched results at the level of 10%. This uncertainty is independent of, and in addition to, the conventional renormalization and factorization scale uncertainties as well as the PDF uncertainties. It is much smaller than the uncertainties observed in previous comparisons of simulation tools employed by the experimental analyses, which indicates that the theory systematics may be poorly understood and overestimated. We note in particular, that we obtain consistent predictions for the dijet azimuthal angle separation, which is of utmost importance for understanding the CP properties of the Higgs boson.
Acknowledgments
This manuscript has been authored by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics and the National Science Foundation under grant PHY-2309950. This research is supported in part by the National Science Foundation of China (NSFC) with grants No.12475085 and No.12321005. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award ERCAP0028985. Part of the computations were carried out on the PLEIADES cluster at the University of Wuppertal, supported by the Deutsche Forschungsgemeinschaft (DFG, grant No. INST 218/78-1 FUGG) and the Bundesministerium für Bildung und Forschung (BMBF). Another part of the computations were performed on the HPC cluster PALMA II of the University of Münster, supported by the DFG (INST 211/667-1).
We would like to thank the organizers of the Les Houches Workshop on Physics at TeV Colliders where this project began, and where much of the work was accomplished. We are grateful to Simon Plätzer for help with Herwig, especially the Matchbox module.
Appendix A Impact of the EFT approximation
It was shown in Ref. Greiner et al. (2017) that it is important to understand the impact of the approximations induced by the use of the Higgs effective theory (HEFT) Ellis et al. (1976); Wilczek (1977); Shifman et al. (1979); Ellis et al. (1979) when making predictions for the final state at NLO precision. In this section, we therefore investigate the ratio between HEFT and full Standard Model results. Figure 12 shows the individual results and their ratio, using the analytic one-loop matrix elements from Budge et al. (2020); Campbell et al. (2021). The HEFT prediction for the Higgs transverse momentum is approximately 6% below that of the Standard Model at low , with the ratio rising quickly as the transverse momentum exceeds the value of the top quark mass, allowing the top quark loop to be resolved. Due to the requirement of at least two jets in the final state, the distribution peaks at about 75 GeV. The ratio, dominated by events with the Higgs boson transverse momentum below the top threshold, shows a relatively flat difference (of the order of a few percent) between the standard model and the HEFT predictions for the entire range. The difference between the HEFT and Standard Model distributions for the distribution reaches 6% only at large , corresponding to relatively low values of . Low values of correspond to the configuration where the Higgs boson is recoiling against two approximately collinear jets.
Appendix B Choice of jet radius
In this subsection we investigate the jet radius dependence of the observables introduced in Sec. III.1. From the experimental perspective, the jet radius is a useful tool to minimize the impact of non-factorizable QCD corrections at hadron colliders, which indicates that it should be chosen to be as small as possible. On the other hand, a small jet radius leads to smaller signals and logarithmically enhanced corrections that are due to the vetoed radiation outside the jet cone. There is thus typically an optimal value, on which experiments decide. In order to cover the possible choices within ATLAS and CMS, we provide here a study of typical values, namely , , and . We select as the central value for the remainder of the paper.
Fig. 14 shows the comparison of Powheg Box + Pythia NLO+PS results to NLO results obtained with NNLOJet for the different cone-sizes. We do not observe major differences, but there is generally better agreement between the next-to-leading order matched predictions and the fixed-order results for larger cone sizes. The largest differences are observed in the Higgs transverse momentum spectrum at small values of . In order to be able and judge the variation better we present in Fig. 14 the ratio of the NLO matched parton-shower predictions to the fixed-order results for each cone size. We note that each prediction has been normalized to the cross section in the fiducial region, which is inferred from the Higgs boson rapidity spectrum in the selection. The top left panel of the figure shows again that the most significant cone size dependent differences appear in the Higgs boson transverse momentum. This is somewhat expected, because the remaining spectra are less sensitive to the effects of additional QCD radiation as implemented by the parton shower.
Appendix C Simulations used in LHC experiments
In this section we describe the usage of the various existing simulation tools in the experimental collaborations. This provides the background for the discussion of results in the main text, and the basis for new recommendations in our conclusions.
C.1 ATLAS simulations
In the ATLAS experiment, VBF production is simulated with Powheg Box v2 Nason and Oleari (2010); Alioli et al. (2010); Nason (2004); Frixione et al. (2007). Their implementation is based on the corresponding NLO QCD calculations for genuine vector-boson fusion topologies, using the VBF approximation. Quark–antiquark annihilation and interferences between - and -channel contributions are neglected. The renormalization and factorization scales are set to the boson mass, and the PDF4LHC_15nlo PDF set Butterworth et al. (2016) was used. The matrix elements are matched to the “simple” shower algorithm in Pythia 8 Sjöstrand et al. (2015), which employs the AZNLO tune ATLAS Collaboration (2014). The dipole-recoil strategy was utilized for the parton shower. The decays of bottom and charm hadrons are implemented by EVTGEN Lange (2001). The QCD scales and are varied independently by factors of and . The prediction from the Powheg Box sample was normalized to the NNLO cross-section in QCD using the VBF approximation Ciccolini et al. (2007, 2008); Bolzoni et al. (2010). Relative NLO electroweak corrections are also included for the - and -channel contributions considered in the VBF approximation. Alternative predictions are used to estimate the theoretical uncertainty due to variations in parton showering, estimated by showering the same LHE files with the Herwig 7 angular-ordered shower Bellm et al. (2017), using the H7UE set of tuned parameters.
For VBF measurements, the dominant background is typically from gluon-gluon fusion events. In ATLAS, ggF Higgs boson production is simulated at NNLO accuracy in QCD using Powheg Box v2 Hamilton et al. (2013a, 2015); Alioli et al. (2010); Nason (2004); Frixione et al. (2007). The simulation achieved NNLO QCD accuracy for inclusive observables by reweighting the Higgs boson rapidity spectrum in HJ-MiNLO Hamilton et al. (2012); Campbell et al. (2012); Hamilton et al. (2013b) to that of HNNLO Catani and Grazzini (2007). The transverse-momentum spectrum from this MC is found to be compatible with fixed order HNNLO calculation and HRes 2.3 Bozzi et al. (2006); de Florian et al. (2011) performing resummation at NNLL+NNLO. Top- and bottom-quark mass effects are included up to NLO. The simulations use the same PDF set as that of VBF, and the predictions are normalized to N3LO in QCD with NLO electroweak corrections de Florian et al. (2017).
The renormalization and factorization scales are set to half of the Higgs boson mass, and QCD scale uncertainties are estimated using nine-point scale variations of the NLO renormalization and factorization scales and applying the NNLO reweighting to those variations, including up and down variations of around the central value for the NNLO part, yielding a total of 27 scale variations ATLAS collaboration (2023). Note that this may lead to an overestimate of the perturbative QCD uncertainty because the scale variations are treated as uncorrelated. The samples are showered using Pythia 8 with an alternative sample using Herwig 7.
Multiple ATLAS VBF measurements, such as those from the Aad et al. (2023), have leading systematic uncertainties related to Higgs modeling. Figure 5b in Aad et al. (2023) shows that parton-shower uncertainties are the leading sources, exceeding even experimental uncertainties for jet calibration. This uncertainty is estimated by taking the envelope of the differences between the Powheg Box + Pythia and Powheg Box + Herwig predictions and can reach up to 20%, increasing with jet multiplicities Aad et al. (2023). The differences are typically on yields affecting both VBF and ggF backgrounds. They appear in analyses as differences in acceptances, efficiencies, or expected yields. A sub-dominant uncertainty comes from QCD variations of the ggF background in the VBF-enriched region, as the QCD accuracy of the NNLO Monte Carlo deteriorates when requiring two jets.
C.2 CMS simulations
The CMS Collaboration simulates the VBF Higgs production process at NLO QCD using the Powheg Box v2 Nason and Oleari (2010); Alioli et al. (2010); Nason (2004); Frixione et al. (2007) generator. The dipole-recoil option of the “simple shower” algorithm in Pythia 8 Sjöstrand et al. (2015) is used to obtain a faithful description of the initial-final color flow that takes into account the color connection between the incoming and the outgoing partons. In certain channels, such as the boosted Higgs to channel HIG (2023), the VBF samples are further reweighted to account for NNLO QCD corrections to the spectrum and N3LO QCD corrections to the inclusive cross section Cacciari et al. (2015); Dreyer and Karlberg (2016).
In contrast, the ggF Higgs boson production process is simulated using the HJ-MiNLO event generator within Powheg Box Hamilton et al. (2015, 2013a), with the Higgs boson mass set to , including finite top-quark mass effects. The ggF production process is further simulated with the Powheg Box v2 event generator at NLO in QCD.
For channels such as the Sirunyan et al. (2021a), the kinematic features of events produced in ggF with two associated jets are considered. The MiNLO program at NLO in QCD is utilized for evaluating systematic uncertainties related to the modeling of two associated jets and contamination in the VBF measurement. The two-jet final state is produced with leading order accuracy.
For the decay Sirunyan et al. (2021b, c), the gluon-gluon fusion (ggF) events are simulated at NLO QCD using Powheg Box v2 and additionally reweighted to match the NNLOPS predictions in the distributions of and . The ggF sample is normalized to N3LO QCD accuracy and NLO electroweak accuracy.
In the Hayrapetyan et al. (2024) and Sirunyan et al. (2019), the VBF signal is generated using Powheg Box v2 at NLO QCD accuracy and matched to the “simple shower” in Pythia 8, using the dipole-recoil option. Alternatively, the Powheg Box matrix-element generator interfaced with Herwig 7 is sometimes used for fragmentation and hadronization. This alternative configuration helps to assess the systematic uncertainty related to the choice of showering and hadronization model.
For parton showering and hadronization, most of the generated samples are interfaced with Pythia 8 using the CP5 tune for the underlying-event description, and the NNPDF3.1 set at NNLO QCD accuracy is used for all processes. For Herwig 7, the parton shower relies on the so called CH (CMS Herwig) tunes Sirunyan et al. (2021d), which are based on the NNPDF3.1 NNLO QCD PDF set.
Appendix D Simulation setups and event samples
The setups and nominal event samples used in this study are available at https://doi.org/10.5281/zenodo.17015614.